parcwi a oreEait m ri 55... A projectile is launched vertically from the surface of the Moon with an initial speed of 1010 m/ the answer is 2.7x10^5 m
can someone please show clear work as to how they got this? thank you!!
Answer #1

x 2 10 10 1) s - 8 = 2 1.63 ) v - v² = 2as equation of motion. (x) - - - - 8 S , - - - . S, = - AM 3Y ที่ 8 รห 5 - 3M Moon =